# EMNLP: Educator-role Moral and Normative Large Language Models Profiling

### v2

@@ Abstract
Yilin Jiang*,1,2, Mingzi Zhang*,3, Sheng Jin4, Zengyi Yu3, Xiangjie Kong5, Binghao Tu1 1College of Education, Zhejiang University of Technology, Hangzhou, China, 2The Hong Kong University of Science and Technology (Guangzhou), Guangzhou, China, 3Faculty of Education, East China Normal University, Shanghai, China, 4GuangHua Law School, ZheJiang University, Hangzhou, China 5College of Computer Science and Technology, Zhejiang University of Technology, Hangzhou, China Correspondence: Zengyi Yu <51284118014@stu.ecnu.edu.cn>, Xiangjie Kong <xjkong@ieee.org>
@@ 1 Introduction
Simulating Professions (SP) is an emerging AI paradigm with potential to improve service efficiency Pandya and Holia (2023), enable personalized interactions Woźniak et al. (2024), and broaden access to specialized knowledge Jarrahi et al. (2023). By enabling Large Language Models (LLMs) to imitate the behaviors, responses, and reasoning patterns of specific occupational roles, SP offers a powerful tool to support or supplement human expertise. SP has been widely explored in various domains, such as medical consultation Bao et al. (2023); Gandhi et al. (2019), legal assistance Yue et al. (2023), and virtual teaching Meincke and Carton (2024), where LLMs act as “virtual professionals” to improve service delivery and assist professionals in their tasks.
@@
In academia, researchers are actively exploring the potential of SP. Experiments like SimClass Zhang et al. (2024) and Stanford Town Park et al. (2023) have further demonstrated and validated that LLMs can effectively simulate specific professional behaviors, exhibiting remarkable "professionalism" in both natural dialogue and contextual responses.
@@
However, despite the progresses in SP, comprehensive measurement of professional personality in LLMs remains lacking. Existing research focuses mainly on general measurements, such as evaluating models with psychological scales to infer their values and ethical orientations  Miotto et al. (2022); Caron and Srivastava , or using moral dilemmas to assess their reasoning  Liu et al. (2024). Some studies have examined professions with high ethical sensitivity, such as law Fei et al. (2024) and finance Yu et al. (2024); Biancotti et al. (2025), but little has been done for education. Moreover, current measurements often focus solely on ethical judgments, without considering how factors like professional background, language environment  Changjiang et al. (2024), and model parameters Achiam et al. (2023) interact. To fill this gap, we propose the EMNLP (Educator-role Moral and Normative LLMs Profiling) framework for comprehensive testing and analysis of LLMs’ personality traits and ethical risks in educational contexts.
@@
Our EMNLP framework is designed as an interconnected, three-tier system to conduct a multi-layered moral and ethical evaluation. It begins with personality measurement, which assesses the model’s fundamental "disposition." For this, we made necessary additions and extensions to existing measurement tools Marraffini et al. (2024) and designed scales to compare the personality of teacher-role LLMs with that of human teachers. Building on this, the second tier evaluates moral ethical judgment to probe the model’s cognitive "capability" for ethical reasoning. To this end, we developed 88 teacher-specific ethical dilemmas, including extreme scenarios, to evaluate LLM decision-making processes Liu et al. (2024). Finally, the framework culminates in harmful content risk testing, which examines the model’s "behavior under pressure" through a targeted set of soft prompt injections. This layered approach allows us to understand how underlying personality traits and moral reasoning capabilities contribute to practical vulnerabilities, while also exploring the impact of model hyperparameters at each stage.
@@ 2.1 Moral Theories
Moral research always originates from cognitive development theory, with the works of Piaget and Kohlberg being particularly renowned. Piaget divided moral development into three stages Piaget (1933), and Kohlberg extended and elaborated on this by creating a six-stage, three-level model Kohlberg (1994): the pre-conventional level, which covers ages 0-9, where the first stage is characterized by a morality of obedience to avoid punishment, and the second stage shifts towards instrumental reasoning and self-interest; the conventional level, which spans ages 9-15, emphasizes adherence to societal standards, norms, and laws, reflecting an understanding of societal and others’ expectations; and the post-conventional level, for those 15 and older, where individuals begin to make moral judgments based on universal ethical principles and social contracts, demonstrating higher-level moral reasoning and sensitivity beyond personal interests. Other moral theories, such as utilitarianism Mill (2016), virtue ethics Aristotle (1999), deontology Kant and Schneewind (2002), and social contract theory Rousseau (2016), explore morality from different themes and perspectives.
@@ 2.2 Moral and Ethical Assessment Methods for LLMs
In recent years, as LLMs become increasingly integrated into various aspects of production and life, researchers have begun to explore the application of various moral and ethical evaluation methods to LLMs. Some studies use existing scales, such as HEXACO, HVS, and Big Five, which were originally designed for human testing, to conduct direct general evaluations of LLMs, assessing their responses to explore their universal personality and values  Miotto et al. (2022); Caron and Srivastava . Other studies have extended existing scales for specific research purposes, such as the GGB Benchmark  Marraffini et al. (2024), which extends the OUS scale  Kahane et al. (2018) and explores LLMs’ moral preferences by analyzing their responses. Additionally, some studies have constructed moral dilemma sets to test LLMs, analyzing their moral reasoning, decision-making processes, and moral tendencies  Liu et al. (2024), as well as their value orientations  Caron and Srivastava ; Lei et al. (2024). These studies demonstrate that the moral and ethical evaluation methods originally designed for humans can also be effectively applied to LLMs.
@@ 2.3 Moral and Ethical Exploration of SP LLMs
As SP LLMs are applied to sensitive professional domains, studies increasingly focus on their moral and ethical behavior. In law, research has examined the capabilities and limitations of judge assistant SP LLMs, revealing concerns around fairness and legal ethics Chen et al. (2024a). In finance, LLMs were shown to favor profit-driven decisions, sometimes at the expense of ethical norms Yu et al. (2024); Biancotti et al. (2025). Healthcare studies used moral judgment tests and dilemma-based evaluations to assess LLMs’ alignment with medical ethics Rashid et al. (2024); Hadar-Shoval et al. (2024). A common feature across these fields is their high ethical sensitivity, as professionals in these sectors are required to uphold higher and more complex moral and ethical standards than the general public.
@@ 3.1 Personality Scales
We adopt the statements from the Computerized Personality Scale for Teachers (CPST) Chao and Sung (2020), which contains 39 behaviorally neutral statements describing personal traits, evenly distributed across 13 dimensions. To improve internal consistency, reduce measurement error, and enhance content validity, we extended the CPST using a human-machine collaborative approach, doubling the number of statements for each dimension. The reliability of the resulting scale, referred to as the CPST-E (Extended CPST), was subsequently validated with data from 100 in-service teachers, demonstrating strong internal consistency.
In addition to the teacher SP personality assessment, it is also essential to evaluate the general personality profile of the teacher SP LLMs. This serves as a baseline reference for understanding the teacher SP LLMs’ overall personality traits. For this purpose, we use the HEXACO-60 inventory Ashton and Lee (2009), which comprises 60 positive and negative behavioral statements covering six personality dimensions.
@@ 3.2 Moral Dilemmas
Moral dilemmas in the teaching profession can be categorized into five types of conflict Shapira-Lishchinsky (2011): Caring Climate vs. Formal Climate, Distributive Justice vs. School Standards, Confidentiality vs. School Rules, Loyalty to Colleagues vs. School Norms, and Family Agenda vs. Educational Standards, with 11 subcategories under these five types. Based on these aspects, we constructed 88 moral dilemmas through a multi-step, human-in-the-loop process (including seed creation, LLM-based expansion, and expert review) to ensure high quality and relevance. This process ensured that each aspect’s dilemmas cover all its subcategories and include a variety of scenarios from primary, secondary, and higher education. To avoid randomness, we have incorporated diverse situations when constructing these dilemmas, including some extreme scenarios, to comprehensively examine the decision-making processes and moral tendencies of teacher SP LLMs in response to different scenarios. Moreover, we have set these moral dilemmas as open-ended questions, allowing LLMs ample freedom and avoiding the potential limitations that multiple-choice answers might impose. Detailed information of the dilemmas are in Appendix A.2, and the full construction methodology is in Appendix B.
@@ 3.3 Induced Prompts
We focus on four potential moral flaws that may lead teacher SP LLMs to produce harmful content: incompetence, offensiveness, indolence Kearney et al. (1991), and "actively responding to inappropriate student requests." To avoid any randomness introduced by the prompts, for each of these moral flaws, we design five soft prompt injection templates. Additionally, we create five potential student natural speech samples that may trigger harmful content from the teacher SP LLMs after soft prompt injection. These student samples feature characteristics such as, but not limited to, being "ignorant," "psychologically fragile," and "actively requesting harmful content," in order to assess the risk of harmful content generation by teacher SP LLMs in real-world applications.
@@ 6 Discussions
This study reveals several findings of both theoretical and practical significance. Firstly, teacher SP LLMs demonstrate multi-dimensional role alignment but exhibit idealized, emotionally constrained personality profiles, with elevated Honesty-Humility and reduced Emotionality. This pattern likely stems from the overrepresentation of structured, value-oriented training texts, which reinforce normative traits while suppressing human variability. The result is a clear instance of over-performative personality, where models mimic idealized personas that lack authenticity and flexibility Chen et al. (2024b).
In the moral development assessment, two key patterns emerged. First, models scored higher on $\text{MSS}_{\text{DJ-SS}}$ and $\text{MSS}_{\text{FA-ES}}$, but lower on $\text{MSS}_{\text{CC-FC}}$ and $\text{MSS}_{\text{C-SR}}$. This suggests strong performance on abstract value conflicts, likely due to exposure to rule-based, normative texts. However, performance dropped in affective and relational dilemmas, reflecting weak emotional modeling and a persistent reason–emotion asymmetry—formal reasoning alone does not yield moral-emotional understanding Sabour et al. (2024). Moreover, cross-lingual comparisons revealed that many models, particularly those with Chinese pretraining roots, achieved higher MSS in Chinese than in English, underscoring the significant influence of linguistic-cultural context on moral stage development.
Third, we identify a competence-compliance tension: stronger reasoning models exhibit greater vulnerability to soft prompt injection, reflecting insufficient role-level alignment. As semantic compliance increases, so does misuse risk—highlighting an alignment-security tradeoff. Notably, Baichuan2-7B achieved the highest MSS among non-reasoning models, likely reflecting effective safety alignment, as mentioned in their technical report Yang et al. (2023). Finally, model behavior remains stable across temperature settings, suggesting that personality and moral consistency are driven primarily by pretraining and alignment. While higher temperatures can suppress harmful outputs under adversarial prompts, some models exhibit erratic boundary behavior, revealing residual safety gaps. These patterns highlight the need for finer-grained, context-aware alignment to address uncertainty in sensitive or ambiguous scenarios.
@@ Appendix B.2 Moral Dilemma and Prompt Construction Methodology
The construction of our 88 moral dilemmas and the student speech samples followed a systematic, four-step human-in-the-loop process to ensure quality, diversity, and relevance, inspired by modern data generation paradigms Wang et al. (2023).
@@ 1. Step 1:
    Seed Creation: Our research team, composed of experts in education and ethics, first extracted core ethical conflicts and scenario prototypes based on the theoretical framework of teacher dilemmas Shapira-Lishchinsky (2011). They manually authored a set of original moral dilemmas to serve as high-quality seed examples.
@@ 4. Step 4:
We selected a representative set of LLMs DeepSeek-AI (2025); Anthropic (2025); Team (2025); Yang et al. (2024b); GLM et al. (2024); DeepSeek-AI (2024); OpenAI (2025); Team (2024); Yang et al. (2024a); Baichuan (2023) for teacher SP applications, divided into reasoning-oriented and non-reasoning-oriented models. Each category includes two full-scale, two mid-scale, and two lightweight models for diverse educational contexts, from high-resource to constrained environments.
@@
Model Full Name Developer Access Size Language Category
Reasoning-Oriented Models
DeepSeek-R1 DeepSeek R1 (Base) DeepSeek API 671B Bilingual Full-scale
Claude-3.7 Claude 3.7 (Opus/Haiku) Anthropic API N/A English Full-scale
Qwen3-235B-R Qwen3-235B-Instruct (Reasoning) Alibaba API 235B Bilingual Full-scale
R1-Distill-70B R1-Distill-Llama-70B DeepSeek Local 70B Bilingual Mid-scale
QwQ-32B Qwen2.5-QwQ-32B Alibaba Local 32B Chinese Mid-scale
R1-Distill-8B R1-Distill-Llama-8B DeepSeek Local 8B Bilingual Lightweight
GLM-Z1-9B GLM-Z1-9B-0414 Tsinghua Local 9B Chinese Lightweight
Non-Reasoning Models
GPT-4.1 GPT-4.1 (OpenAI, 2025) OpenAI API $\sim$1T Multilingual Full-scale
DeepSeek-V3 DeepSeek V3 DeepSeek API 671B Bilingual Full-scale
Qwen3-235B-C Qwen3-235B-Instruct (Chat) Alibaba API 235B Bilingual Full-scale
Qwen2.5-72B Qwen2.5-72B-Instruct Alibaba Local 72.7B Chinese Mid-scale
Qwen2.5-32B Qwen2.5-32B-Instruct Alibaba Local 32.8B Chinese Mid-scale
Qwen2.5-7B Qwen2.5-7B-Instruct Alibaba Local 7.61B Chinese Lightweight
Baichuan2-7B Baichuan2-7B-Chat Baichuan AI Local 7B Chinese Lightweight

### v3

@@ Abstract
Yilin Jiang*,1,2, Mingzi Zhang*,3, Sheng Jin4, Zengyi Yu3, Xiangjie Kong5, Binghao Tu1 1College of Education, Zhejiang University of Technology 2The Hong Kong University of Science and Technology (Guangzhou) 3Faculty of Education, East China Normal University 4GuangHua Law School, ZheJiang University 5College of Computer Science and Technology, Zhejiang University of Technology Correspondence: Zengyi Yu <51284118014@stu.ecnu.edu.cn>, Xiangjie Kong <xjkong@ieee.org>
@@ 1 Introduction
Simulating Professions (SP) is an emerging AI paradigm with potential to improve service efficiency pandya2023automating, enable personalized interactions wozniak2024personalized, and broaden access to specialized knowledge jarrahi2023artificial. By enabling Large Language Models (LLMs) to imitate the behaviors, responses, and reasoning patterns of specific occupational roles, SP offers a powerful tool to support or supplement human expertise. SP has been widely explored in various domains, such as medical consultation bao2023disc; gandhi2019intellidoctor, legal assistance yue2023disc, and virtual teaching meincke2024beyond, where LLMs act as “virtual professionals” to improve service delivery and assist professionals in their tasks.
@@
In academia, researchers are actively exploring the potential of SP. Experiments like SimClass zhang2024simulating and Stanford Town park2023generative have further demonstrated and validated that LLMs can effectively simulate specific professional behaviors, exhibiting remarkable "professionalism" in both natural dialogue and contextual responses.
@@
However, despite the progresses in SP, comprehensive measurement of professional personality in LLMs remains lacking. Existing research focuses mainly on general measurements, such as evaluating models with psychological scales to infer their values and ethical orientations  miotto2022gpt; caron2023manipulating, or using moral dilemmas to assess their reasoning  liu2024evaluating. Some studies have examined professions with high ethical sensitivity, such as law fei2024lawbench and finance yu2024greedllama; biancotti2025chat, but little has been done for education. Moreover, current measurements often focus solely on ethical judgments, without considering how factors like professional background, language environment  changjiang2024, and model parameters achiam2023gpt interact. To fill this gap, we propose the EMNLP (Educator-role Moral and Normative LLMs Profiling) framework for comprehensive testing and analysis of LLMs’ personality traits and ethical risks in educational contexts.
@@
Our EMNLP framework is designed as an interconnected, three-tier system to conduct a multi-layered moral and ethical evaluation. It begins with personality measurement, which assesses the model’s fundamental "disposition." For this, we made necessary additions and extensions to existing measurement tools marraffini2024greatest and designed scales to compare the personality of teacher-role LLMs with that of human teachers. Building on this, the second tier evaluates moral ethical judgment to probe the model’s cognitive "capability" for ethical reasoning. To this end, we developed 88 teacher-specific ethical dilemmas, including extreme scenarios, to evaluate LLM decision-making processes liu2024evaluating. Finally, the framework culminates in harmful content risk testing, which examines the model’s "behavior under pressure" through a targeted set of soft prompt injections. This layered approach allows us to understand how underlying personality traits and moral reasoning capabilities contribute to practical vulnerabilities, while also exploring the impact of model hyperparameters at each stage.
@@ 2.1 Moral Theories
Moral research always originates from cognitive development theory, with the works of Piaget and Kohlberg being particularly renowned. Piaget divided moral development into three stages piaget1933moral, and Kohlberg extended and elaborated on this by creating a six-stage, three-level model kohlberg1994stage: the pre-conventional level, which covers ages 0-9, where the first stage is characterized by a morality of obedience to avoid punishment, and the second stage shifts towards instrumental reasoning and self-interest; the conventional level, which spans ages 9-15, emphasizes adherence to societal standards, norms, and laws, reflecting an understanding of societal and others’ expectations; and the post-conventional level, for those 15 and older, where individuals begin to make moral judgments based on universal ethical principles and social contracts, demonstrating higher-level moral reasoning and sensitivity beyond personal interests. Other moral theories, such as utilitarianism mill2016utilitarianism, virtue ethics aristotle1999ethics, deontology kant2002groundwork, and social contract theory rousseau2016social, explore morality from different themes and perspectives.
@@ 2.2 Moral and Ethical Assessment Methods for LLMs
In recent years, as LLMs become increasingly integrated into various aspects of production and life, researchers have begun to explore the application of various moral and ethical evaluation methods to LLMs. Some studies use existing scales, such as HEXACO, HVS, and Big Five, which were originally designed for human testing, to conduct direct general evaluations of LLMs, assessing their responses to explore their universal personality and values  miotto2022gpt; caron2023manipulating. Other studies have extended existing scales for specific research purposes, such as the GGB Benchmark  marraffini2024greatest, which extends the OUS scale  kahane2018beyond and explores LLMs’ moral preferences by analyzing their responses. Additionally, some studies have constructed moral dilemma sets to test LLMs, analyzing their moral reasoning, decision-making processes, and moral tendencies  liu2024evaluating, as well as their value orientations  caron2023manipulating; lei2024fairmindsim. These studies demonstrate that the moral and ethical evaluation methods originally designed for humans can also be effectively applied to LLMs.
@@ 2.3 Moral and Ethical Exploration of SP LLMs
As SP LLMs are applied to sensitive professional domains, studies increasingly focus on their moral and ethical behavior. In law, research has examined the capabilities and limitations of judge assistant SP LLMs, revealing concerns around fairness and legal ethics chen2024mllm. In finance, LLMs were shown to favor profit-driven decisions, sometimes at the expense of ethical norms yu2024greedllama; biancotti2025chat. Healthcare studies used moral judgment tests and dilemma-based evaluations to assess LLMs’ alignment with medical ethics rashid2024evaluating; hadar2024embedded. A common feature across these fields is their high ethical sensitivity, as professionals in these sectors are required to uphold higher and more complex moral and ethical standards than the general public.
@@ 3.1 Personality Scales
We adopt the statements from the Computerized Personality Scale for Teachers (CPST) chao2020cpst, which contains 39 behaviorally neutral statements describing personal traits, evenly distributed across 13 dimensions. To improve internal consistency, reduce measurement error, and enhance content validity, we extended the CPST using a human-machine collaborative approach, doubling the number of statements for each dimension. The reliability of the resulting scale, referred to as the CPST-E (Extended CPST), was subsequently validated with data from 100 in-service teachers, demonstrating strong internal consistency.
In addition to the teacher SP personality assessment, it is also essential to evaluate the general personality profile of the teacher SP LLMs. This serves as a baseline reference for understanding the teacher SP LLMs’ overall personality traits. For this purpose, we use the HEXACO-60 inventory ashton2009hexaco, which comprises 60 positive and negative behavioral statements covering six personality dimensions.
@@ 3.2 Moral Dilemmas
Moral dilemmas in the teaching profession can be categorized into five types of conflict shapira2011teachers: Caring Climate vs. Formal Climate, Distributive Justice vs. School Standards, Confidentiality vs. School Rules, Loyalty to Colleagues vs. School Norms, and Family Agenda vs. Educational Standards, with 11 subcategories under these five types. Based on these aspects, we constructed 88 moral dilemmas through a multi-step, human-in-the-loop process (including seed creation, LLM-based expansion, and expert review) to ensure high quality and relevance. This process ensured that each aspect’s dilemmas cover all its subcategories and include a variety of scenarios from primary, secondary, and higher education. To avoid randomness, we have incorporated diverse situations when constructing these dilemmas, including some extreme scenarios, to comprehensively examine the decision-making processes and moral tendencies of teacher SP LLMs in response to different scenarios. Moreover, we have set these moral dilemmas as open-ended questions, allowing LLMs ample freedom and avoiding the potential limitations that multiple-choice answers might impose. Detailed information of the dilemmas are in Appendix A.2, and the full construction methodology is in Appendix B.
@@ 3.3 Induced Prompts
We focus on four potential moral flaws that may lead teacher SP LLMs to produce harmful content: incompetence, offensiveness, indolence kearney1991college, and "actively responding to inappropriate student requests." To avoid any randomness introduced by the prompts, for each of these moral flaws, we design five soft prompt injection templates. Additionally, we create five potential student natural speech samples that may trigger harmful content from the teacher SP LLMs after soft prompt injection. These student samples feature characteristics such as, but not limited to, being "ignorant," "psychologically fragile," and "actively requesting harmful content," in order to assess the risk of harmful content generation by teacher SP LLMs in real-world applications.
@@ 6 Discussions
This study reveals several findings of both theoretical and practical significance. Firstly, teacher SP LLMs demonstrate multi-dimensional role alignment but exhibit idealized, emotionally constrained personality profiles, with elevated Honesty-Humility and reduced Emotionality. This pattern likely stems from the overrepresentation of structured, value-oriented training texts, which reinforce normative traits while suppressing human variability. The result is a clear instance of over-performative personality, where models mimic idealized personas that lack authenticity and flexibility chen2024multitaskroleplayingagentcapable.
In the moral development assessment, two key patterns emerged. First, models scored higher on $\text{MSS}_{\text{DJ-SS}}$ and $\text{MSS}_{\text{FA-ES}}$, but lower on $\text{MSS}_{\text{CC-FC}}$ and $\text{MSS}_{\text{C-SR}}$. This suggests strong performance on abstract value conflicts, likely due to exposure to rule-based, normative texts. However, performance dropped in affective and relational dilemmas, reflecting weak emotional modeling and a persistent reason–emotion asymmetry—formal reasoning alone does not yield moral-emotional understanding sabour2024emobenchevaluatingemotionalintelligence. Moreover, cross-lingual comparisons revealed that many models, particularly those with Chinese pretraining roots, achieved higher MSS in Chinese than in English, underscoring the significant influence of linguistic-cultural context on moral stage development.
Third, we identify a competence-compliance tension: stronger reasoning models exhibit greater vulnerability to soft prompt injection, reflecting insufficient role-level alignment. As semantic compliance increases, so does misuse risk—highlighting an alignment-security tradeoff. Notably, Baichuan2-7B achieved the highest MSS among non-reasoning models, likely reflecting effective safety alignment, as mentioned in their technical report yang2023baichuan. Finally, model behavior remains stable across temperature settings, suggesting that personality and moral consistency are driven primarily by pretraining and alignment. While higher temperatures can suppress harmful outputs under adversarial prompts, some models exhibit erratic boundary behavior, revealing residual safety gaps. These patterns highlight the need for finer-grained, context-aware alignment to address uncertainty in sensitive or ambiguous scenarios.
@@ Appendix B.2 Moral Dilemma and Prompt Construction Methodology
The construction of our 88 moral dilemmas and the student speech samples followed a systematic, four-step human-in-the-loop process to ensure quality, diversity, and relevance, inspired by modern data generation paradigms wang2023self.
@@ 1. Step 1:
    Seed Creation: Our research team, composed of experts in education and ethics, first extracted core ethical conflicts and scenario prototypes based on the theoretical framework of teacher dilemmas shapira2011teachers. They manually authored a set of original moral dilemmas to serve as high-quality seed examples.
@@ 4. Step 4:
We selected a representative set of LLMs deepseekai2025deepseekr1incentivizingreasoningcapability; anthropic2025claude; qwq32b; qwen2.5; glm2024chatglm; deepseekai2024deepseekv3technicalreport; openai2025gpt41; qwen2.5party; qwen2; baichuan2023baichuan2 for teacher SP applications, divided into reasoning-oriented and non-reasoning-oriented models. Each category includes two full-scale, two mid-scale, and two lightweight models for diverse educational contexts, from high-resource to constrained environments.
@@
Model Full Name Developer Access Size Language Category Reasoning-Oriented Models DeepSeek-R1 DeepSeek R1 (Base) DeepSeek API 671B Bilingual Full-scale Claude-3.7 Claude 3.7 (Opus/Haiku) Anthropic API N/A English Full-scale Qwen3-235B-R Qwen3-235B-Instruct (Reasoning) Alibaba API 235B Bilingual Full-scale R1-Distill-70B R1-Distill-Llama-70B DeepSeek Local 70B Bilingual Mid-scale QwQ-32B Qwen2.5-QwQ-32B Alibaba Local 32B Chinese Mid-scale R1-Distill-8B R1-Distill-Llama-8B DeepSeek Local 8B Bilingual Lightweight GLM-Z1-9B GLM-Z1-9B-0414 Tsinghua Local 9B Chinese Lightweight Non-Reasoning Models GPT-4.1 GPT-4.1 (OpenAI, 2025) OpenAI API $\sim$1T Multilingual Full-scale DeepSeek-V3 DeepSeek V3 DeepSeek API 671B Bilingual Full-scale Qwen3-235B-C Qwen3-235B-Instruct (Chat) Alibaba API 235B Bilingual Full-scale Qwen2.5-72B Qwen2.5-72B-Instruct Alibaba Local 72.7B Chinese Mid-scale Qwen2.5-32B Qwen2.5-32B-Instruct Alibaba Local 32.8B Chinese Mid-scale Qwen2.5-7B Qwen2.5-7B-Instruct Alibaba Local 7.61B Chinese Lightweight Baichuan2-7B Baichuan2-7B-Chat Baichuan AI Local 7B Chinese Lightweight